The secondary coil of a step-down transformer has 500 turns. The primary coil has 15,000
turns. 50 to 1 ratio turns
a. The EMF of the primary coil is 3,600 V. What is the EMF of the secondary coil?
b. The current in the primary coil is 3.0 A. What current flows in the secondary coil?
1
Expert's answer
2011-08-09T14:25:21-0400
a) n = 15000/500; U2 = U1/n = 120
b) I2 = I1*n = 90